For the reaction: B2O3 + 3H2O 3O2 +B2H6 H=2035KJ

How much energy is needed to react excess B2O3 with 25.0 g of H2O
Answer #1

Answer:- Energy needed = 941.6 kJ

Solution:-

B2O3(s) + 3H2O(g) ---> B2H6(g) + 3O2(g) ΔH = 2035 kJ

This means 2035 kJ of energy is needed to react excess B2O3 with 25.0 g of H2O

Lets convert 25.0 g of H2O into moles

Number of moles = weight/molecular weight = 25/ 18.01 = 1.38812 moles

1 mole of B2O3 reacts with 3 mole of H2O

so that

0.46271 mole of B2O3 reacts with 1.38812 mole of H2O

Energy needed = 0.46271 * 2035

Energy needed = 941.6 kJ